solve the initial value problem x_dx^dy-2y=〖2x〗^4, y(2)=8
1
Expert's answer
2020-12-02T01:51:40-0500

"xy'-2y=(2x)^4"

Solution of homogenуous equation we will find in the form "y=Ax^k" :

"kAx^k-2Ax^k=0"

"k=0"

"y_{h}=Ax^2"

Solution of inhomogeneous equation we will find as "y=B(x)x^2" :

"y'=B'(x)x^2+2B(x)x"

After substitution into equation:

"B'(x)x^3+2B'(x)x^2-2B(x)x^2=(2x)^4"

"B'(x)=16x"

"B(x)=8x^2+C"

"y=(8x^2+C)x^2"

"y(2)=(8*4+C)*4=8"

"C=2-32=-30"


Answer: "y=8x^4-30x^2"